Depreciation Methods

Gruman Company purchased a machine for $198,000 on January 2, 2016. It made the following estimates:

Service Life 5 years or 10,000 hours
Production 180,000 units
Residual value $18,000

In 2016, Gruman uses the machine for 1,700 hours and produces 45,000 units. In 2017, Gruman uses the machine for 1,400 hours and produces 30,000 units. If required, round your final answers to the nearest dollar.

Required:

1. Compute the depreciation for 2016 and 2017 under each of the following methods:

Straight-line method

2016 $
2017 $

Sum-of-the-years'-digits method

2016 $
2017 $

Double-declining-balance method

2016 $
2017 $

Activity method based on hours worked

2016 $
2017 $

Activity method based on units of output

2016 $
2017 $

2. For each method, what is the book value of the machine at the end of 2016? At the end of 2017?

Straight-line method

2016 $
2017 $

Sum-of-the-years'-digits method

2016 $
2017 $

Double-declining-balance method

2016 $
2017 $

Activity method based on hours worked

2016 $
2017 $

Activity method based on units of output

2016 $
2017 $

3. If Gruman used a service life of 8 years or 15,000 hours and a residual value of $9,000, what would be the effect on the following under the straight-line, sum-of-the-years'-digits, and double-declining-balance depreciation methods?

Depreciation expense

Straight-line method

2016 $
2017 $

Sum-of-the-years'-digits method

2016 $
2017 $

Double-declining-balance method

2016 $
2017 $

Book value

Straight-line method

2016 $
2017 $

Sum-of-the-years'-digits method

2016 $
2017 $

Double-declining-balance method

2016 $
2017 $
Answer #1

1. A) Depreciation under Straight Line Method

   Under Staight Line Method Cost of Asset is written off equally during it's useful.         

    Hence depreciation for each useful life will be equal.

      Depreciation = (Purchase Cost of Machine - Salvage Vale) / Service Life

      Depreciation = (198000 - 18000) / 5

                              = 180000 / 5 = 36000 $/ Year

     Depreciation for 2016 = 36000 $

    Depreciation for 2017 = 36000 $

B) Depreciation under Sum of Years Digit Method

This method takes assets expected life and adds together the digits of each year. Then each years digit is diveded by this sum to determine percentage by which the Asset shoud be depeciated each year starting with the highest number in Year 1.

Sum of Useful Life = 5+4+3+2+1 = 15

Depreciation percentage for 2016 = 5/15 * 100 = 33.33 %

Depreciation percentage for 2017 = 4/15 * 100 = 26.67 %

Depreciation for 2016 = (198000 - 18000) * 33.33 %

                                        = 180000 * 33.33 % = 60000 $

Depreciation for 2017 = (198000 - 18000) * 26.67 %

                                        = 180000 * 26.67 % = 48000 $

C) Depreciation under Double Declining balance method

Under this method Asset depreciated twice at the rate which is depreciated under Straight Line Method and each years depreciation is deducted from book value to calculate depreciation in the following year. Depreciation is calculated till it reaches it's Salvage Value

Depreciation = 2 * Cost of Asset * Depreciation Rate OR

                            2 * Cost of Asset / Service Life

Rate of Depreciation = 1 / Service Life * 100

                                       = 1 / 5 * 100 = 20 %

Depreciation for 2016 = 2 * 198000 * 20 %

                                        = 396000 * 20 % = 79200 $

Depreciation for 2017 = (198000 - 79200) * 2 * 20%

                                       = 118800 * 2 * 20 % = 47520 $

D)    Activity method based on Hours Worked

   Under this method Depreciation is calaculated based on hours worked each year to Total useful working hours of Machine.

Depreciation = Hours worked / Total Useful Hours * Cost of Asset

Depreciation for 2016 = 1700 / 10000 * (198000 -18000)           

                                        = 180000 * 1700 /10000

                                       = 30600 $

Depreciation for 2017 = 1400 / 10000 * (198000 -18000)      

=180000 * 1400 /10000

   = 25200 $

E) Activity method based on Units of Output

Under this method Depreciation is calaculated based on Units produced each year to Total Units Capacity.

Depreciation = Cost of Asset * Units Produced / Total Units Capacity

Depreciation for 2016 = (198000 -18000) * 45000 /180000

                                        = 180000 * 45000 /180000 = 45000 $

Depreciation for 2016 = (198000 -18000) * 30000 / 180000

                                        = 180000 * 30000 / 180000 = 30000 $

2) A. Book Value under Straight Line Method

Year Cost / WDV (A) Depreciation(B) Book Value (A - B)
2016 180000 36000 144000
2017 144000 36000 108000

B. Book Value under Sum of Years Digit Method

Year Cost / WDV (A) Depreciation(B) Book Value (A - B)
2016 180000 60000 120000
2017 120000 48000 72000

C. Book Value under Double Declining balance method

Year Cost / WDV (A) Depreciation(B) Book Value (A - B)
2016 198000 79200 118800
2017 118800 47520 71280

D. Book Value under Activity method based on Hours Worked

Year Cost / WDV (A) Depreciation(B) Book Value (A - B)
2016 180000 30600 149400
2017 149400 25200 124200

E. Book Value under Activity method based on Units of Output

Year Cost / WDV (A) Depreciation(B) Book Value (A - B)
2016 180000 45000 135000
2017 135000 30000 105000

3) Effects of Changes

Service Life = 8 Years / 15000 Hours

Residual Value = 9000 $

A) Depreciation under Straight Line Method

   Depreciation = 198000 - 9000 / 8

                           = 189000 / 8 = 23625 $ (for every year)

B) Depreciation under Sum of Years Digit Method

Sum of Useful Life = 8+7+6+5+4+3+2+1 = 36

Depreciation percentage for 2016 = 8 / 36 * 100 = 22.22 %

Depreciation percentage for 2017 = 7 / 36 * 100 = 19.44 %

Depreciation for 2016 = 189000 * 22.22 % = 42000 $

Depreciation for 2017 = 189000 * 19.44 % = 36750 $

C) Depreciation under Double Declining balance method

Rate of Depreciation = 1 / Service Life * 100

                                       =1/8 *100 = 12.5 %

Depreciation for 2016 = 2 * 198000 * 12.5 %

                                        = 396000 * 12.5 % = 49500 $

Depreciation for 2017 = (198000 - 49500) * 2 * 12.5%

                                       = 148500 * 2 * 12.5 % = 37125 $

D) Book Value under Straight Line Method

Year Cost / WDV (A) Depreciation(B) Book Value (A - B)
2016 189000 23625 165375
2017 165375 23625 141750

E) Book Value under Sum of Years Digit Method

Year Cost / WDV (A) Depreciation(B) Book Value (A - B)
2016 189000 42000 147000
2017 147000 36750 110250

F) Book Value under Double Declining balance method

Year Cost / WDV (A) Depreciation(B) Book Value (A - B)
2016 198000 49500 148500
2017 148500 37125 111375
  • Change in Capacity in Hours only effects calaculation of Depreciation under Activity method based on Hours Worked